RC meeting and Run Plan, Monday December 20th

From SBS wiki
Jump to: navigation, search

RC Meeting Information

Run Coordinator: Bogdan Wojtsekhowski

Location: Counting House Conference Room

Time: 3:15 pm

Remote connection link: https://bluejeans.com/256205934

Toll-Free Number (U.S.& Canada): 1-888-240-2560

Bluejeans Meeting ID: 256 205 934

Upcoming Shifts Which Need Filled

VERY URGENT: Shifts in January need help - 195 openings to fill

Today's estimate: Optic beam Tue. Jan 11, "Bonus beam" Jan 12-16. Official beam Jan 17

Accelerator and Hall Status

  • Preparation for beam off period
  • End of beam: Tue. Dec. 21 at 7:00 AM
  • Tuesday 12/21 at 8pm expected test of the Access Control System for one hour

To be Done During Extended Down Time

  • Replace J0 INFN plane with UVa plane UV#4 and related electronics
  • Spectrometers move to positions for SBS-14 kinematics
  • Cosmic tests of HCal and BBcal
  • Study of GRINCH
  • Checks quality of the fiber connections within DAQ

Run Plan

  • Continue data taking on LD2 for "Cycle 24"

Opportunistic Access Completed

  • None

Resolved Issues

  • Oscilloscope on SL monitor for instant check of the delays
  • GEM sector loss => reduced HV by 25 V on all GEMs
  • HCAL 12 channels HV drop

Unresolved Issues

  • NEW: HCAL ROC17 ERROR for last 12+ hours
  • Occasional timing shifts of ~70 ns
  • GEM efficiency analysis.
    • Use of the Y-tgt cut for determination of the actual efficient of the detector.

8 uA => 11 uA

Subsystem Updates

  1. DAQ general
    1. Alex update. Statistics of DAQ reboots.
  2. HCal
    1. F1 errors (HCal ROC17 sync errors)... Don't ignore!! End, Reset, Configure, Download. If that doesn't work, call DAQ on call.
    2. Could 50k replay show a problem FLAG?
    3. News on time resolution and amplitude analysis.
  3. BBCal
    1. Lesson from BBcal HV loss for few hours on Friday. Update on monitoring/reboot methods.
  4. Timing Hodoscope
    1. Update on TH setup and time resolution. Plan for spare parts.
  5. GRINCH
    1. Update from data analysis.
  6. GEMS
    1. DAQ update
    2. HV and efficiency status. Perspectives for higher luminosity.
  7. Beamline
    1. Update on raster analysis.
  8. Replay analysis
    1. Massive data replay: script preparation is in progress.
    2. Observations on data quality.